OpenStudy (anonymous):

Math Analysis: A fan spins 100 times in 1 minute. Through how many degrees does a point on the tip of the wheel move in 15 seconds?

Parth (parthkohli):

15 seconds is 1/4 of a minute. Then it'd spin 25 times because 25 is 1/4 of 100. One turn is 360 degrees so: \( \color{Black}{\Rightarrow 25 \times 360 }\)

OpenStudy (cwtan):

1minute =60seconds 60=15*4 so In 15seconds it turn for 100/4=25 Angle turned= 25*360

OpenStudy (anonymous):

isn't there a formula for this :P